if cos25 0 + sin25 0 =p, then cos50 0 is (A) ∫ 2-p 2 (b) – ∫ 2-9 2 © p ∫ 2-p 2 (d) – p ∫ 2-p 2 this symbol represent ∫ square root . pl. show the working

if cos250 + sin250 =p, then cos500 is
(A) ∫ 2-p2          (b) – ∫ 2-92     © p ∫ 2-p2  (d)  – p ∫ 2-p2
 
this symbol represent ∫  square root . pl. show the working

Grade:11

1 Answers

Vikas TU
4 years ago
Dear student
Cos25 + Sin25 = K .........1 
Cos2Q = Cos^2Q - Sin^Q 
=> (CosQ+SinQ)(CosQ-SinQ)
Square equation 1 
Cos^25 +Sin^2 25 + 2Cos25*Sin25 = K^2
1+ sin80 = k^2
Sin 80 = k^2 -1 
Draw traingle and apply pythagorus theorem 
Cos 50 = ksqrt(2-k^2)